OpenStudy (anonymous):

PLZ HELP! MEDAL REWARD!!:) On a given planet, the weight of an object varies directly with the mass of the object. Suppose that an object whose mass is 3kg weighs 18N. Calculate the mass of another object that weighs 24N.

OpenStudy (mathstudent55):

Use a proportion.

OpenStudy (anonymous):

i was never thought that

OpenStudy (imstuck):

Use the direct variation proportion y = kx. Do you know that?

OpenStudy (mathstudent55):

A proportion is imply two ratios set equal to each other. The ratio of weight to mass is always the same. For the given 3 kg mass that weight 18 N, what is the ratio of weight to mass?

OpenStudy (imstuck):

Not proportion...equation.

OpenStudy (mathstudent55):

If you want to do it with an equation, you can also do that. Let's use the variables m for mass and w for weight.

OpenStudy (anonymous):

You have to use an proportion, Mathstudent is correct. @IMStuck

OpenStudy (mathstudent55):

w = km Plug in the known values of 3 for mass and 18 for weight. What do you get for k? 18 = k * 3 k = 18/3 = 6 Now we know k = 6, so the equation is: w = 6m Now plug in 24 for w, and solve for m.

OpenStudy (anonymous):

m =4kg..?

OpenStudy (mathstudent55):

With a proportion, this is how you do it: \(\dfrac{18~N}{3~kg} = \dfrac{24~N}{x} \) Cross multiply: 18x = 24 * 3 18x = 72 Solve for x.

OpenStudy (anonymous):

4

OpenStudy (mathstudent55):

Correct.